# Mut4All: Fuzzing Compilers via LLM-Synthesized Mutators Learned from Bug Reports

### v1

@@ 3. Approach
Figure 3. The overview of Mut4All.
We now describe the design of Mut4All, with a detailed explanation of the processes underlying each step shown in the examples in Section 2. Figure 3 shows an overview of Mut4All, briefly describing the workflow and detailing the inputs and outputs of each component. Overall, Mut4All takes as input historical bug reports, prompt templates for each LLM agent, and existing manually written mutator examples, and produces a set of validated mutators ready for integration into our subsequent compiler fuzzing phase. At a finer granularity, as mentioned above, Mut4All comprises three specialized LLM agents corresponding to three key steps: (1) the Mutator Invention agent, which takes bug reports and bug-triggering code snippets as input and outputs a set of mutator specifications; (2) the Mutator Implementation Synthesis agent, which takes these specifications, along with our prompt templates and human-written mutator examples, to implement raw mutators in the target language; and (3) the Mutator Refinement agent, which takes the raw mutators and their associated compiler error messages (if any) as input and outputs validated mutators ready for large-scale compiler fuzzing. We now describe the design of each agent in detail.
@@ 3.1. Mutator Invention
Compilers, as the foundation of modern software development, are among the most actively maintained and scrutinized components in the open-source ecosystem. Bugs and usage issues are frequently reported on platforms such as GitHub, often accompanied by detailed discussions, triggering code snippets, and diagnostic outputs. These bug reports reflect valuable domain knowledge from compiler developers and often highlight specific language constructs or internal behaviors prone to failure. Mut4All leverages these historical bug reports as input to its first component, the Mutator Invention agent. This agent analyzes the bug descriptions and associated code to identify underlying language features responsible for the reported failures, and uses this analysis to design targeted mutation operators. It is based on the observation that many historical compiler bugs arise from recurring constructs or subtle corner cases that are difficult to expose via generic mutations—such as simple arithmetic rewrites or control-flow transformations. By grounding the mutation logic in real-world failure scenarios, Mut4All can generate semantics-aware, feature-specific mutators that are more likely to trigger previously unseen bugs.
@@ 3.2. Mutator Implementation Synthesis
Once a mutator has been invented, the next step is to synthesize its executable implementation. This typically involves writing AST transformation code that modifies a program according to the intended mutation—a process that is tedious and error-prone when done manually, especially for complex and evolving compiler infrastructures. To address this, Mut4All introduces a second agent responsible for Mutator Implementation Synthesis. This LLM-based agent generates transformation code tailored to the target language based on the mutator specification. To ensure code quality and semantic correctness, we fine-tune the LLM using OpenAI’s standard fine-tuning API, guided by 10 human-written mutator examples per language. These examples cover diverse syntactic constructs and representative transformation patterns, serving as domain-specific demonstrations that help the model synthesize accurate and idiomatic AST manipulation code.
To enforce structural consistency, Mut4All adopts a predefined mutator template for each supported language. The template defines the skeleton that the LLM must complete with appropriate mutation logic like matching AST nodes, applying transformations, and handling edge cases. Listing 4 and Listing 5 present simplified examples for Rust and C++, respectively. Each template specifies the mutator’s entry point, filtering conditions, and placeholders for inserting LLM-generated transformation code.
@@
10 fn chain_of_thought(&self) -> &str {
11 ""
12 }
13}
Listing 4: The core part of the mutator template for Rust.
@@
3 if (!MT || !Result.Context->getSourceManager().isWrittenInMainFile(MT->getLocation()))
@@
6}
Listing 5: The core part of the mutator template for C++.
@@ 3.3. Mutator Refinement
If one mutator fails to compile, or if applying it to any seed program causes a compilation error, Mut4All captures the full compiler error messages (e.g., due to missing imports, incorrect type handling, or invalid AST rewrites). These messages are embedded into a structured prompt template (Figure 6) that instructs the LLM to revise the mutator while preserving its intended transformation semantics. The refinement process is conducted in an iterative loop. After each repair attempt, the revised mutator is recompiled and re-applied across all seed programs. If the mutator passes compilation and does not cause any errors when applied to the test suite, it is marked as valid and added to the mutator pool for fuzzing. Otherwise, the process continues for up to a maximum of $N$ iterations (we use $N=10$ in our experiments). If the mutator still fails validation after the limit is reached, it is discarded.
Furthermore, to ensure that the mutator performs meaningful transformations, we include a post-validation check: if the mutator compiles and runs but does not introduce any syntactic change in any test input (e.g., due to overly strict matching logic), it is considered ineffective and also discarded. This ensures that the final mutator set is not only valid, but actively contributes to provide meaningful and diverse mutations for the fuzzing phase.
@@
3.4. Fuzzing Loop
After collecting a set of validated mutators, Mut4All enters its core fuzzing loop to systematically test target compilers. Algorithm 1 outlines this process in detail.
Given an initial set of seed programs $S$, Mut4All first expands the seed corpus using rule-based transformations via the enhance function, following the fixed-replacement strategy introduced by Clozemaster (Gao et al., 2025), which replaces each identifier in the seed programs with a randomly chosen identifier of the same type. This step aims to diversify the seed space and uncover additional transformation opportunities. The enhanced seed corpus is denoted as $\mathcal{P}$. In each fuzzing iteration, Mut4All randomly selects $i$ mutators from the validated mutator set $M$ and applies them to each program in the current seed pool $\mathcal{P}$. The resulting mutated programs $p^{\prime}$ are then compiled and executed on each target compiler $c$ in the compiler set $C$. If a compiler crashes or hangs, the corresponding input is immediately recorded as a bug-triggering example. For successfully compiled programs, the execution results $r$ of the generated binaries are collected. Mut4All compares these outputs across compilers or against expected values (when available). If any inconsistency is detected—such as divergent outputs, abnormal termination, or unexpected side effects—a behavioral bug is reported. The fuzzing loop continues until a predefined timeout $T$ is reached or until other stopping criteria (e.g., maximum number of iterations or detected bugs) are met.
@@ 3.4. Fuzzing Loop
4 while $\operatorname{time\_not\_expired}$ ($T$) do
@@
6     $M^{\prime}\leftarrow\operatorname{random\_select\_N\_mutators}$ ($M$, $i$)
@@
9           $p^{\prime}\leftarrow\operatorname{apply}(p,m)$
10
@@ 11          $R\leftarrow\phi$
13              $\operatorname{res,exe}\leftarrow\operatorname{compile}$($c$, $p$)
14              if $\operatorname{crash(res)}~OR~\operatorname{hang(res)}$  then
[…]
16
17             else if $\{\operatorname{exe}\}\neq\phi$ then
18                 $r\leftarrow\operatorname{execute}$($\operatorname{exe}$)
@@ 19                 $R\leftarrow R\cup\{\operatorname{res}\}$
20
21
22          if $\operatorname{has\_inconsistency}$ ($R$) then
23              $\operatorname{record}$ ($p$, $R$)
24
25
@@ 23              $\operatorname{record}$ ($p$, $R$)
Algorithm 1 The fuzzing loop of Mut4All.
@@ 3. RQ3
    How effective does Mut4All in mutator synthesis for compiler fuzzing?
@@ 4.1. Implementation
To demonstrate the cross-language capability of Mut4All, we select two prominent system-level languages, Rust and C++, as a proof of concept. We implement a dedicated web crawler to extract fixed or resolved bug reports from the official GitHub repositories [^5] https://github.com/rust-lang/rust and https://github.com/llvm/llvm-project of rustc and Clang, and parse their contents to extract the embedded bug-triggering programs. For each target language, we construct the initial seed corpus by combining all extracted bug-triggering programs with the official test suites of the corresponding compilers (Zhong, 2022). This process yields 20,527 seeds for Rust and 27,786 seeds for C++. We expand the corpus to 100,000 programs using Clozemaster-style rule-based mutations (Gao et al., 2025), which replace each identifier in the seed programs with a randomly selected identifier of the same type to diversify the seed space. We then discard programs that fail to compile, resulting in 87,733 Rust and 68,176 C++ programs. Additionally, from the crawled bug reports, we randomly select 500 reports per language for the Mutator Invention agent to synthesize language-specific mutators. In total, Mut4All generates 319 valid mutators for Rust and 403 valid mutators for C++.
We adopt GPT-4o (Achiam et al., 2023) as the underlying LLM, which also offers an API for custom fine-tuning. We utilize this capability to further improve the quality and correctness of the generated mutators by providing 10 manually crafted examples that cover a diverse range of AST transformations as additional fine-tuning data for each language.
For the bug-hunting experiments, we allocate a fuzzing time budget of 48 hours per language and initiate the fuzzing loop using all valid mutators. Within this time budget, we observe that the fuzzing pipelines for both Rust and C++ typically remain in the stage of applying up to two mutators per seed.

### v2

@@ 3. Approach
Figure 2. The overview of the interaction of the three Mut4All agents.
We now describe the design of Mut4All, with a detailed explanation of the processes underlying each step shown in the examples in Section 2. Figure 2 shows an overview of Mut4All agents, briefly describing the workflow and detailing the inputs and outputs of each component. Overall, Mut4All takes as input historical bug reports, prompt templates for each LLM agent, and existing manually written mutator examples, and produces a set of validated mutators ready for integration into our subsequent compiler fuzzing phase. At a finer granularity, as mentioned above, Mut4All comprises three specialized LLM agents corresponding to three key steps: (1) the Mutator Invention agent, which takes bug reports and bug-triggering code snippets as input and outputs a set of mutator specifications; (2) the Mutator Implementation Synthesis agent, which takes these specifications, along with our prompt templates and human-written mutator examples, to implement raw mutators in the target language; and (3) the Mutator Refinement agent, which takes the raw mutators and their associated compiler error messages (if any) as input and outputs validated mutators ready for large-scale compiler fuzzing. We now describe the design of each agent in detail.
@@ 3.1. Mutator Invention
Compilers, as the foundation of modern software development, are among the most actively maintained and scrutinized components in the open-source ecosystem. Bugs and usage issues are frequently reported on platforms such as GitHub, often accompanied by detailed discussions, triggering code snippets, and diagnostic outputs. These bug reports reflect valuable domain knowledge from compiler developers and often highlight specific language constructs or internal behaviors prone to failure. Mut4All leverages these historical bug reports as input to its first component, the Mutator Invention agent. This agent analyzes the bug descriptions and associated code to identify underlying language features responsible for the reported failures, and uses this analysis to design targeted mutation operators. It is based on the observation that many historical compiler bugs arise from recurring constructs or subtle corner cases that are difficult to expose via mutators that are supported by existing approaches. By grounding the mutation logic in real-world failure scenarios, Mut4All can generate semantics-aware, feature-specific mutators that are more likely to trigger previously unseen bugs.
@@ 3.2. Mutator Implementation Synthesis
Once a mutator has been invented, the next step is to synthesize an executable implementation. This typically involves writing AST transformation code that modifies a program according to the intended mutation—a process that is tedious and error-prone when done manually, especially for complex and evolving compiler infrastructures. To address this, Mut4All introduces a second agent responsible for Mutator Implementation Synthesis. This LLM-based agent generates transformation code tailored to the target language based on the mutator specification. To ensure code quality and semantic correctness, we fine-tune the LLM using OpenAI’s standard fine-tuning API, guided by 10 human-written mutator examples per language. These examples cover diverse syntactic constructs and representative transformation patterns, serving as domain-specific demonstrations that help the model synthesize accurate and idiomatic AST manipulation code.
To enforce structural consistency, Mut4All adopts a predefined mutator template for each supported language. The template defines the skeleton that the LLM must complete with appropriate mutation logic like matching AST nodes, applying transformations, and handling edge cases. Listing 5 and Listing 6 present simplified examples for Rust and C++, respectively. Each template specifies the entry point of one mutator, filtering conditions, and placeholders for inserting LLM-generated transformation code. These mutator templates are general and expressive enough to capture the full operational logic of all mutators, providing a unified structure for matching and transforming code.
@@
12}
Listing 5: The core part of the mutator template for Rust.
1// Mutating the target AST node (i.e., insertion, deletion, replacement)
@@
4 if (!MT $|$ !Result.Context->getSourceManager().isWrittenInMainFile(MT->getLocation()))
@@
7}
8// Localizing the target AST node, and binding the mutation operations
[…]
10 MatchFinder matchFinder;
11 DeclarationMatcher matcher = {{MatchRule}}.bind({{NodeName}});
12 Callback callback(TheRewriter);
13 matchFinder.addMatcher(matcher, &callback);
14 matchFinder.matchAST(Context);
15}
Listing 6: The core part of the mutator template for C++.
@@ 3.3. Mutator Refinement
If one mutator fails to compile, or if applying it to any seed program causes a compilation error, Mut4All captures the full compiler error messages (e.g., due to missing imports, incorrect type handling, or invalid AST rewrites). These messages are embedded into a structured prompt template (Figure 5) that instructs the LLM to revise the mutator while preserving its intended transformation. The refinement process is conducted in an iterative loop. After each repair attempt, the revised mutator is recompiled and re-applied across all seed programs. If the mutator implementation successfully compiles and does not cause any errors when applied to the test suite, it is marked as valid and added to the mutator pool for fuzzing. Otherwise, the process continues for up to a maximum of $N$ iterations (we use $N=10$ in our experiments). If the mutator still fails validation after the limit is reached, it is discarded.
Furthermore, to ensure that the mutator performs meaningful transformations, we include an additional check: if the mutator compiles and runs but does not introduce any syntactic change in any test input (e.g., due to overly strict matching logic), it is considered ineffective and also discarded. This ensures that the final mutator set is not only valid, but actively contributes to provide meaningful and diverse mutations for the fuzzing phase.
@@
3.4. Fine-Tuning
Due to the frequent evolution of compiler infrastructures and their underlying AST manipulation libraries, many APIs encoded in the pretrained LLM’s knowledge have become outdated or deprecated across compiler versions, which can lead to incorrect or obsolete API usage when synthesizing mutators. To mitigate this issue, we fine-tune the Mutator Implementation Synthesis agent on a set of up-to-date, manually designed mutator specifications, enabling it to learn the canonical and version-consistent use of modern AST APIs. Specifically, we manually construct 10 mutator specifications for each language as demonstration data. For Rust, we construct 10 mutator specifications and their corresponding implementations derived from representative bug reports; for C++, we build upon 10 mutators from our prior work, TyMut (Wang et al., 2025b), and refine both their specifications and implementations based on real bug reports. These carefully designed mutator–specification pairs jointly serve as supervised examples for fine-tuning, covering diverse AST manipulation patterns and language-specific features observed in real compiler ecosystems.
We fine-tune GPT using Supervised Fine-Tuning (SFT), where a token-level cross-entropy loss based on maximum likelihood estimation is applied to minimize the discrepancy between generated text and human demonstrations. The loss function is defined as:
(1) $\mathcal{L}_{\text{SFT}}(\theta)=-\sum_{t=1}^{T}\log P_{\theta}(y_{t}\mid y_{<t},x)$
where $x$ denotes the input prompt, $y=(y_{1},y_{2},\dots,y_{T})$ is the target sequence of human demonstrations with length $T$. $P_{\theta}(\cdot)$ represents the conditional probability distribution parameterized by the model parameters $\theta$, and $y_{<t}$ denotes the prefix of tokens before the current time step $t$. In our case, the human demonstrations correspond to manually constructed mutator–specification pairs, where each specification serves as the input prompt and its corresponding mutator implementation serves as the target output. By minimizing this loss, the model is trained to maximize the likelihood of producing the human-written target sequence, thus aligning its generation behavior with up-to-date, version-stable compiler APIs.
3.5. The Main Fuzzing Procedure
@@
5
6while $\neg\textnormal{{Expired(}}\textnormal{\emph{T}}\textnormal{{)}}$ do
7
@@
9     $M^{\prime}\leftarrow\textnormal{{RandomPick(}}\textnormal{\emph{M, i}}\textnormal{{)}}$
@@
12           $p^{\prime}\leftarrow\textnormal{{Apply(}}\textnormal{\emph{p,m}}\textnormal{{)}}$
@@ 13           $R\leftarrow\phi$
15              ${res,exe}\leftarrow\textnormal{{Compile(}}\textnormal{\emph{c, p'}}\textnormal{{)}}$
16              if $\textnormal{{Crash(}}\textnormal{\emph{res}}\textnormal{{)}}~\operatorname{OR}~\textnormal{{Hang(}}\textnormal{\emph{res}}\textnormal{{)}}$ then
17                 Record(p, c)
18
19             else if $\{{exe}\}\neq\phi$ then
20                 $r\leftarrow\textnormal{{Execute(}}\textnormal{\emph{exe}}\textnormal{{)}}$
@@ 21                 $R\leftarrow R\cup\{\operatorname{res}\}$
22
23
24          if Inconsistent(R) then
25              Record(p, R)
@@
29
30
Algorithm 1 The Fuzzing Procedure of Mut4All.
After collecting a set of validated mutators, Mut4All enters its core fuzzing loop to test target compilers systematically. Algorithm 1 outlines this process in detail.
Given an initial set of seed programs $S$, Mut4All first expands the seed corpus using rule-based transformations via the enhance function (introduced in detail in Section 3.6). This step aims to diversify the seed space and uncover additional transformation opportunities, which will be elaborated on later. The enhanced seed corpus is denoted as $\mathcal{P}$. During the fuzzing loop, in each iteration, Mut4All randomly selects $i$ mutators from the validated mutator set $M$ and applies them to each program in the current seed pool $\mathcal{P}$. The resulting mutated programs $p^{\prime}$ are then compiled and executed on each target compiler $c$ in the compiler set $C$. If a compiler crashes or hangs, the corresponding input is immediately recorded as a bug-triggering example. For successfully compiled programs, the execution results $r$ of the generated binaries are collected. Mut4All compares these outputs across compilers or against expected values (when available). If any inconsistency is detected—such as divergent outputs, abnormal termination, or unexpected side effects—a behavioral bug is recorded. The fuzzing loop continues until a predefined timeout $T$ is reached or until other stopping criteria (e.g., maximum number of iterations or detected bugs) are met. For all recorded programs, we manually inspected them for potential undefined behavior and cross-checked existing bug reports to determine whether they are duplicates before submitting new issues to the open-source compiler communities.
3.6. Adaptive Search-Based Seed Enhancement
Input: Seed set of programs $\mathcal{P}$
Input: Retry limit $T$
Input: The lower bound of weight $w_{lb}\in[0,1]$
Input: Decay factor $\gamma\in(0,1)$
Output: Mutated seed program set $\mathcal{P}^{\prime}$
1
2
$\mathcal{P}^{\prime}\leftarrow\mathcal{P}$
// initialize the output set
3
$\mathsf{Index}\leftarrow\emptyset$
// $\mathsf{Index}:\text{AST type}\mapsto\{\text{subtrees}\}$
4
5foreach $\mathsf{prog}\in P$ do
6  foreach $n$ in $\mathsf{prog}$ do
[…]
     $w(n)\leftarrow w_{lb}+(1-w_{lb})/$Len($n$)
     // initialize weights
8
9
10
111ex
$\mathsf{Used}\leftarrow\emptyset$
// record selected donor subtrees
12
13foreach $\mathsf{prog}\in P$ do
14  $\widehat{\mathsf{prog}}\leftarrow\mathsf{prog}$
15  while preorder traversing nodes $n$ of $\widehat{\mathsf{prog}}$ do
16     $\mathsf{choose}\leftarrow$ Random $(0,1)$
17     if $\mathsf{choose}<w(n)$ then
18       continue
    $S\leftarrow\{d~|~d\in\mathsf{Index}[\textnormal{{TypeOf(}}\textnormal{\emph{n}}\textnormal{{)}}]\land w(d)>\mathsf{choose}\}$
     // filter target donor nodes
19     if $S=\emptyset$ then
20       Continue
21
22    for $t\leftarrow 1$ to $T$ do
        $d\leftarrow\textsc{UniformChoice}(S)$
        // pick donor node
23
24       if $d\in\mathsf{Used}$ then
[…]
26       $\widehat{\mathsf{prog}}^{\prime}\leftarrow\textnormal{{Replace}}(\widehat{\mathsf{prog}},n,d)$
27        if Compiles($\widehat{\mathsf{prog}}^{\prime}$) then
28           $\mathcal{P}^{\prime}\leftarrow\mathcal{P}^{\prime}\cup\{\widehat{\mathsf{prog}}\}$
29           $\mathsf{Used}\leftarrow\mathsf{Used}\cup\{d\}$
30       else
31           $w(d)\leftarrow w_{lb}$
32
33
34
35
36return $\mathcal{P}^{\prime}$
Algorithm 2 Adaptive Seed Enhancement.
To diversify the seed programs and amplify the effectiveness of compiler fuzzing (Wen et al., 2023; Wu et al., 2023c; Gao et al., 2024b), we design an adaptive subtree-level mutation algorithm (i.e., one that operates on entire AST subtrees rather than individual nodes or tokens) that dynamically prioritizes promising mutation locations based on feedback from compilation success. Unlike conventional token- or statement-level mutation strategies (Wen et al., 2023; Wu et al., 2023c), our subtree-level design preserves syntactic validity by ensuring that all replacements occur between type-compatible AST fragments, thereby maintaining compilation feasibility while exploring semantically diverse variants. Algorithm 2 outlines this procedure.
Given an initial seed set $\mathcal{P}$, the algorithm first builds an AST node type–indexed subtree pool, mapping each AST node type to all subtrees of that type found in $\mathcal{P}$, and assigns each node an initial mutation weight (Lines 1–6). This index serves as a repository of mutation candidates for type-preserving replacements. The weight of each node is determined by its subtree size, ensuring that smaller subtrees are more likely to be selected while bounding the minimum weight with a user-defined lower limit $w_{lb}$ (Line 6).
For each seed program (Lines 8–26), the algorithm traverses its AST in preorder (Lines 10–26). At each node $n$, it probabilistically decides whether to attempt a mutation based on its weight $w(n)$ (Lines 11–13). If selected, the algorithm retrieves all subtrees of the same AST node type as $n$ from the index whose weights exceed the randomly sampled threshold (Line 14). If no suitable candidates exist, the current node $n$ is skipped (Line 16). Otherwise, the algorithm attempts up to $T$ replacements (Line 17). In each attempt, it randomly samples a donor subtree $d$ from the candidate set (Line 18). If $d$ has been used previously, its weight is decayed (Line 20). The node $n$ is then replaced with $d$, and the mutated program is compiled (Line 22). The algorithm enforces type compatibility between donor and target subtrees based on their AST node kinds and inferred types.
If compilation succeeds, the mutated program is added to the output set $\mathcal{P}^{\prime}$, and $d$ is marked as used (Lines 23–24). Otherwise, the weight of $d$ is reset to the lower bound (Line 26). To account for the possibility that $d$ may still be valid in other code contexts where $n$ appears, we maintain a non-zero lower bound for all weights. Finally, the enhanced seed set $\mathcal{P}^{\prime}$ is returned, providing a richer and more structurally diverse corpus for the subsequent fuzzing campaign introduced in Section 3.5.
@@ 3. RQ3
    What are the costs of using Mut4All for mutator synthesis?
  4. RQ4
    What is the effectiveness of the mutators synthesized by Mut4All?
@@ 4.1. Implementation
To demonstrate the cross-language capability of Mut4All, we select two prominent system-level languages, Rust and C++, as a proof of concept. We implement a dedicated web crawler to extract fixed or resolved bug reports from the official GitHub repositories of rustc [^5] https://github.com/rust-lang/rust and Clang [^6] https://github.com/llvm/llvm-project, and parse their contents to extract the embedded bug-triggering programs. Both repositories are from production compilers and well-maintained open-source communities.
For each target language, we construct the initial seed corpus by combining all extracted bug-triggering programs with the official test suites of the corresponding compilers (Zhong, 2022). This process yields 20,481 seeds for Rust and 27,786 seeds for C++. We expand the corpus by our adaptive seed enhancement approach (i.e., Algorithm 2) for Rust and C++. In our implementation, we set $\gamma=0.95$, $w_{lb}=0.3$, and $T=200$. The decay factor $\gamma$ ensures that a node’s selection weight decays to the lower bound after approximately 30 unsuccessful attempts, while the non-zero lower bound $w_{lb}=0.3$ guarantees that every node still retains a small probability of being selected. Finally, we collect the expanded seed programs, which consist of 87,688 Rust and 68,176 C++ programs, respectively. Additionally, from the crawled bug reports, we randomly select 500 reports per language for the Mutator Invention agent to synthesize language-specific mutators. In total, Mut4All generates 319 valid mutators for Rust and 403 valid mutators for C++.
We adopt GPT-4o (Achiam et al., 2023) as the underlying LLM, which also offers an API for custom fine-tuning. We utilize this capability to further improve the quality and correctness of the LLM-synthesized mutators. For fine-tuning, we curated a set of 20 manually written mutators—10 for Rust and 10 for C++—to provide the LLM with high-quality examples of AST-level transformations and correct API usage. On average, each Rust mutator contains 65 LOC and invokes 16 AST manipulation APIs, while each C++ mutator has 86 LOC and 26 API calls. These mutators provide the LLM with a diverse range of high-quality, language-specific demonstrations for learning accurate AST transformation patterns. We adopted the default supervised fine-tuning (i.e., SFT) pipeline provided by OpenAI [^7] https://openai.com/index/gpt-4o-fine-tuning/. The total economic cost of fine-tuning across all stages was approximately 11.8\$. Moreover, we configure different temperature values for the agents to balance creativity and determinism: the Mutator Invention Agent uses a high temperature (0.8) to encourage diverse and exploratory mutation ideas, while the Mutator Implementation Synthesis Agent and Mutator Refinement Agent adopt lower temperatures (0.4 and 0.2, respectively) to ensure more stable and consistent code generation.
For the bug-hunting experiments, we initiate the fuzzing loop using all valid mutators. Within this time budget, we observe that the fuzzing pipelines for both Rust and C++ typically remain in the stage of applying up to two mutators per seed.